OpenStudy (anonymous):

The length of the line segment whose endpoints are (3,2) and (-2,1) is

jimthompson5910 (jim_thompson5910):

use the distance formula d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((-2-3)^2+(1-2)^2) d = sqrt((-5)^2+(-1)^2) I'll let you finish